OpenStudy (anonymous):

Differentiate y=sin^-1(e^3x)

OpenStudy (anonymous):

The derivative of arcsin or sin^-1(x) is dx / sqrt(1 - x^2). In this case x = e^3x and dx = 3e^3x. This yields: y'= [3e^(3x)] / [sqrt( 1 - (e^3x))]